# Homotopy Continuation Made Easy: Regression-based Online Simulation of Starting Problem-Solution Pairs

### v1

@@ 7.4 Real World Experiments
Table 3 shows the median results over all three sequences of LaMAR dataset. We adopted a similar setting for RANSAC as the above section with $500$ maximum iterations. The proposed solver still presents consistent performances, where at least $60\%$ pairs have rotation errors smaller than $2$ degrees together with small translation angular errors on each sequence.
@@ 8 Conclusion
The present paper introduces a new paradigm for correspondence-based geometric problem solving using an elegant three-stage combination of a regression network, an online correspondence simulator, and homotopy continuation. The regression network acts as a general solution approximator and can be trained in simulation only. While pure regression accuracy is often insufficient on its own, a key message conveyed through our results is that the accuracy is generally still good enough to simulate a consistent starting problem-solution pair that successfully enables single-solution continuation. As demonstrated on two generalized camera problems, the strategy enables high success rates on common computing hardware, and the resulting solvers prove to be a viable solution if utilized as part of a random sample and consensus scheme. In particular, we propose a successful solution to the generalized relative pose and scale problem, which is at least 20×20$\times$20 × faster than traditional polynomial elimination techniques.
@@
This supplementary presents the results did not shown in the main paper for the lack of space.

### v2

@@ 7.4 Real World Experiments
Table 3 shows the median results over all three sequences of LaMAR dataset. We adopted a similar setting for RANSAC as the above section with $500$ maximum iterations. The proposed solver still presents consistent performances, where simulator HC has higher success rate with faster running time than Gröbner Bases.
@@ 8 Conclusion
The present paper introduces a new paradigm for correspondence-based geometric problem solving using an elegant three-stage combination of a regression network, an online correspondence simulator, and homotopy continuation. The regression network acts as a general solution approximator and can be trained in simulation only. While pure regression accuracy is often insufficient on its own, a key message conveyed through our results is that the accuracy is generally still good enough to simulate a consistent starting problem-solution pair that successfully enables single-solution continuation. As demonstrated on two generalized camera problems, the strategy enables high success rates on common computing hardware, and the resulting solvers prove to be a viable solution if utilized as part of a random sample and consensus scheme. In particular, we propose a successful solution to the generalized relative pose and scale problem, which is around 17×17$\times$17 × faster than traditional polynomial elimination techniques.
9 Acknowledgments
The authors would like to acknowledge the funding support provided by projects 22DZ1201900, 22ZR1441300, and DFYJBJ-1 by the Natural Science Foundation of Shanghai.
@@ 9 Acknowledgments
This supplementary presents the results not shown in the main paper for the lack of space.
@@ 11 GRPS Noise Resilience Full Results
12 Evaluation within RANSAC
We further compare the proposed $8$pts simulator HC and Gröbner Bases solver inside RANSAC using synthetic data. The following table presents translation error and scale error in percentage.
Table 4: RANSAC experiments for the GRPS problem using Gröbner Bases $\mid$ simulator HC. Image points are randomly perturbed by 2-pixel uniform noise. We use vanilla RANSAC with at most $200$ iterations. As can be observed, outlier ratios of up to $40\%$ are easily handled within 200 iterations. And regressor-based simulator HC is consistently faster than Gröbner Bases.
Ratio Succ. (%) $\mathcal{E}_{\mathbf{R}}(deg)$ $\mathcal{E}_{\mathbf{t}}(\%)$ $\mathcal{E}_{\mathbf{s}}(\%)$ Iters Time(s)
10%   99.5 $\mid$ 100 0.10 $\mid$ 0.09 2.19 $\mid$ 2.03 2.13 $\mid$ 1.98   93 $\mid$ 125 2.33 $\mid$ 0.85
20%   99.4 $\mid$ 100 0.11 $\mid$ 0.10 2.31 $\mid$ 2.11 2.40 $\mid$ 2.23 171 $\mid$ 193 4.37 $\mid$ 1.31
30%   99.4 $\mid$ 100 0.13 $\mid$ 0.14 2.75 $\mid$ 3.40 2.61 $\mid$ 3.17 200 $\mid$ 200 5.10 $\mid$ 1.35
40% 98.9 $\mid$ 91 0.21 $\mid$ 0.29 5.26 $\mid$ 7.57 5.07 $\mid$ 7.81 200 $\mid$ 200 4.59 $\mid$ 1.35
13 More Discussion
We propose a general method to solve nonlinear equations, particularly polynomial systems that challenge traditional approaches such as Gröbner Bases (GB) and Levenberg-Marquardt (LM). Given the limitations of classical methods, researchers have actively explored modern alternatives, such as homotopy continuation (HC) and learning-based approaches. However, these methods either induce high computational cost (HC exhaustively tracks all roots) or suffer from poor performance. Simulator HC significantly enhances the performance of simple learning-based methods while avoiding the computational expenses of tracking all solution paths, as in standard HC.
The proposed Simulator HC could be applied to many applications. Examples of important problems for which no alternative exists and where our approach could be applied are given by: Rolling Shutter Camera Relative Pose Problem (no minimal solver available), Inverse Kinematics Problem (high degrees of freedom). Applying simulator HC to these unsolved problems is a very interesting direction for future research.